# Similarity to contraction semigroups:

### v1

@@ 8.2. Similarity to isometries in terms of operator means
There are several classical results characterizing when a linear operator $E$ with $\sigma(E)\subseteq\mathbb{{R}}$ is similar to a self-adjoint operator, i.e. when $iE$ generates a uniformly bounded $C_{0}$-group. Some of these results are formulated in terms of two-sided bounds for means of $L^{2}$-norms of semigroup orbits, see for instance [56, Section 2, p. 43], [60, Remark 3, p. 15], [78, Theorem 3.1]. However, unlike the approach in the preceding section, they do not employ operator weights.
@@
see also Theorem A.1. Apart from their theoretical importance, these results appeared to be useful in the study of concrete differential operators, see, for example, [45, 57] and the references therein. Our next result extends these criteria to the semigroup setting, providing two-sided bounds for orbit means with operator weights, in the spirit of Theorems 8.1 and 8.6, cf. [46, Proposition 1.15].
@@
If the operator $C$ is closed, then the equivalence (ii) $\iff$ (iii) in Proposition 8.8 follows by standard arguments involving Plancherel’s theorem and two-sided inequalities between Cesàro means and Abel means; see, for instance, [78, Lemma 1.1].
@@ 8.3. Observability and controllability via similarity
In the setting of Theorem 8.11, Grabowski and Callier showed in their seminal work [34] that infinite time exact observability of $(E,C)$ (with unbounded $C$) is equivalent to similarity of $\mathcal{T}$ to a contraction semigroup, assuming that $\mathcal{T}$ is exponentially stable (see Theorem $3.1$ and Corollary $3.1$ in [34]). Their argument is rather intricate, as they aim at showing the dissipativity for the generator of $\mathcal{S}$ (given by (8.3)) rather than establishing upper and lower norm bounds for orbits of $\mathcal{S}$, which is considerably more direct. Furthermore, [35, Proposition 5.1] shows that infinite time exact observability implies similarity to a contraction semigroup, though without addressing strong stability. Among other related results in the literature, we mention [53, Proposition 2] where one proves that exponential stability of $\mathcal{T}$, combined with the exact controllability of $(E,I)$, implies similarity of $\mathcal{S}$ to a semigroup of contractions.
@@
Versions of Theorem 8.12 for bounded $C$ where obtained in [23, Theorem 1.3] and [81, Theorem 1]. However, most of PDE applications require $C$ unbounded, and $C$ could be even non-closable even if $(C,E)$ is observable/controllable (as in Remark 8.3). If $C$ is bounded, then, as proved in [81, Theorem 1], the finite time observability of $(C,E)$ is equivalent to the boundedness from below (equivalently, left-invertibility) of $\mathcal{T}$ and thus, by [7, Theorem 7.3], to the extendability of $\mathcal{T}$ to a $C_{0}$-group into a larger Hilbert space $K\supseteq H.$ Thus, in view of Corollary 1.3, the difference between the cases of bounded and unbounded $C$’s in Theorem 8.12 can be interpreted as the difference between the properties of $\mathcal{T}$ to be dilatable and to be extendable to a $C_{0}$-group on larger Hilbert space containing $H$.
@@
The results above provide a different proof of Rota’s type theorem for quasi-contraction $C_{0}$-semigroups, [63, Proposition 5.2], stating that every exponentially stable, quasi-contraction $C_{0}$-semigroup belongs to $\mathcal{SC}(H)$. Indeed, let $\mathcal{T}$ be an exponentially stable $C_{0}$-semigroup in $\mathcal{SQC}(H)$. By Theorem 8.12, there exists a finite time admissible observation operator $C$ for $\mathcal{T}$ such that the pair $(E,C)$ is finite time exactly observable. As $\mathcal{T}$ is exponentially stable, $C$ is an infinite time admissible observation operator for $\mathcal{T}$ making $(E,C)$ exactly observable in infinite time; see, for instance, [77, Proposition 6.5.2]. It then follows from Theorem 8.11 that $\mathcal{T}\in\mathcal{SC}(H)$, as required.
@@ 9.3. Similarity by complete boundedness
It is well known that a bounded operator $T$ on a Hilbert space $H$ is similar to a contraction if and only if the homomorphism $p\to p(T)$ defined on the algebra $\mathcal{P}$ of polynomials $p$ extends to a completely bounded map from the disc algebra $A(\mathbb{D})$ into $\mathcal{L}(H)$; see [67, Theorem 9.11] or [71, Theorem 4.13]. The corresponding result in the context of $C_{0}$-semigroups follows as a direct corollary and is essentially contained in the proof of [27, Theorem 5.1]; see also [48, Theorem 2.2]. Since this criterion plays a fundamental role in the discrete theory, we formulate explicitly its semigroup analogue.
@@
The following result characterizes $\mathcal{T}\in\mathcal{SC}(H)$ in terms of the complete boundedness of $\Theta_{\mathcal{T}}$, thus invoking a well-known concept in operator space theory, see [67, Chapters 8 & 9] and [71, Chapter 3] for precise definitions and background.
@@ Appendix A Similarity to isometric semigroups
Here we prove a generalization of a result stated by Naboko in [60, Section 3] for generators of bounded $C_{0}$-groups (i.e. operators similar to skew-adjoint ones). Such a generalization is needed in the proof of Theorem 8.8, and since the proof in [60] is omitted, we provide a complete argument in the case of isometric $C_{0}$-semigroups and with more general assumptions on resolvent two-sided bounds.
@@
The inequality (A.5) with $n=1$ implies that E-E is a sectorial operator, so it is densely defined; see, for instance, [36, Proposition 2.1.1]. In view of (A.5) $E$ satisfies the assumptions of Hille-Yosida’s theorem and is thus generates a bounded $C_{0}$-semigroup $\mathcal{T}=(T(t))_{t\geq 0}$ on $H$.
@@ Appendix B Basics on observability and controllability
It is well-known (see e.g. [77, p. 145 & Proposition 6.5.2]) that if $\mathcal{T}$ is exponentially stable, then infinite time observability and finite time observability are equivalent. More precisely,
@@
For the proof of this result see e.g. [77, Theorems 4.4.3 & 11.2.1].

### v2

@@ 8.2. Similarity to isometries in terms of operator means
There are several classical results characterizing when a linear operator $E$ with $\sigma(E)\subseteq\mathbb{{R}}$ is similar to a self-adjoint operator, i.e. when $iE$ generates a uniformly bounded $C_{0}$-group. Some of these results are formulated in terms of two-sided bounds for means of $L^{2}$-norms of semigroup orbits, see for instance [57, Section 2, p. 43], [61, Remark 3, p. 15], [79, Theorem 3.1]. However, unlike the approach in the preceding section, they do not employ operator weights.
@@
see also Theorem A.1. Apart from their theoretical importance, these results appeared to be useful in the study of concrete differential operators, see, for example, [46, 58] and the references therein. Our next result extends these criteria to the semigroup setting, providing two-sided bounds for orbit means with operator weights, in the spirit of Theorems 8.1 and 8.6, cf. [47, Proposition 1.15].
@@
If the operator $C$ is closed, then the equivalence (ii) $\iff$ (iii) in Proposition 8.8 follows by standard arguments involving Plancherel’s theorem and two-sided inequalities between Cesàro means and Abel means; see, for instance, [79, Lemma 1.1].
@@ 8.3. Observability and controllability via similarity
In the setting of Theorem 8.11, Grabowski and Callier showed in their seminal work [35] that infinite time exact observability of $(E,C)$ (with unbounded $C$) is equivalent to similarity of $\mathcal{T}$ to a contraction semigroup, assuming that $\mathcal{T}$ is exponentially stable (see Theorem $3.1$ and Corollary $3.1$ in [35]). Their argument is rather intricate, as they aim at showing the dissipativity for the generator of $\mathcal{S}$ (given by (8.3)) rather than establishing upper and lower norm bounds for orbits of $\mathcal{S}$, which is considerably more direct. Furthermore, [36, Proposition 5.1] shows that infinite time exact observability implies similarity to a contraction semigroup, though without addressing strong stability. Among other related results in the literature, we mention [54, Proposition 2] where one proves that exponential stability of $\mathcal{T}$, combined with the exact controllability of $(E,I)$, implies similarity of $\mathcal{S}$ to a semigroup of contractions.
@@
Versions of Theorem 8.12 for bounded $C$ where obtained in [23, Theorem 1.3] and [82, Theorem 1]. However, most of PDE applications require $C$ unbounded, and $C$ could be even non-closable even if $(C,E)$ is observable/controllable (as in Remark 8.3). If $C$ is bounded, then, as proved in [82, Theorem 1], the finite time observability of $(C,E)$ is equivalent to the boundedness from below (equivalently, left-invertibility) of $\mathcal{T}$ and thus, by [7, Theorem 7.3], to the extendability of $\mathcal{T}$ to a $C_{0}$-group into a larger Hilbert space $K\supseteq H.$ Thus, in view of Corollary 1.3, the difference between the cases of bounded and unbounded $C$’s in Theorem 8.12 can be interpreted as the difference between the properties of $\mathcal{T}$ to be dilatable and to be extendable to a $C_{0}$-group on larger Hilbert space containing $H$.
@@
The results above provide a different proof of Rota’s type theorem for quasi-contraction $C_{0}$-semigroups, [64, Proposition 5.2], stating that every exponentially stable, quasi-contraction $C_{0}$-semigroup belongs to $\mathcal{SC}(H)$. Indeed, let $\mathcal{T}$ be an exponentially stable $C_{0}$-semigroup in $\mathcal{SQC}(H)$. By Theorem 8.12, there exists a finite time admissible observation operator $C$ for $\mathcal{T}$ such that the pair $(E,C)$ is finite time exactly observable. As $\mathcal{T}$ is exponentially stable, $C$ is an infinite time admissible observation operator for $\mathcal{T}$ making $(E,C)$ exactly observable in infinite time; see, for instance, [78, Proposition 6.5.2]. It then follows from Theorem 8.11 that $\mathcal{T}\in\mathcal{SC}(H)$, as required.
@@ 9.3. Similarity by complete boundedness
It is well known that a bounded operator $T$ on a Hilbert space $H$ is similar to a contraction if and only if the homomorphism $p\to p(T)$ defined on the algebra $\mathcal{P}$ of polynomials $p$ extends to a completely bounded map from the disc algebra $A(\mathbb{D})$ into $\mathcal{L}(H)$; see [68, Theorem 9.11] or [72, Theorem 4.13]. The corresponding result in the context of $C_{0}$-semigroups follows as a direct corollary and is essentially contained in the proof of [28, Theorem 5.1]; see also [49, Theorem 2.2]. Since this criterion plays a fundamental role in the discrete theory, we formulate explicitly its semigroup analogue.
@@
The following result characterizes $\mathcal{T}\in\mathcal{SC}(H)$ in terms of the complete boundedness of $\Theta_{\mathcal{T}}$, thus invoking a well-known concept in operator space theory, see [68, Chapters 8 & 9] and [72, Chapter 3] for precise definitions and background.
@@ Appendix A Similarity to isometric semigroups
Here we prove a generalization of a result stated by Naboko in [61, Section 3] for generators of bounded $C_{0}$-groups (i.e. operators similar to skew-adjoint ones). Such a generalization is needed in the proof of Theorem 8.8, and since the proof in [61] is omitted, we provide a complete argument in the case of isometric $C_{0}$-semigroups and with more general assumptions on resolvent two-sided bounds.
@@
The inequality (A.5) with $n=1$ implies that E-E is a sectorial operator, so it is densely defined; see, for instance, [37, Proposition 2.1.1]. In view of (A.5) $E$ satisfies the assumptions of Hille-Yosida’s theorem and is thus generates a bounded $C_{0}$-semigroup $\mathcal{T}=(T(t))_{t\geq 0}$ on $H$.
@@ Appendix B Basics on observability and controllability
It is well-known (see e.g. [78, p. 145 & Proposition 6.5.2]) that if $\mathcal{T}$ is exponentially stable, then infinite time observability and finite time observability are equivalent. More precisely,
@@
For the proof of this result see e.g. [78, Theorems 4.4.3 & 11.2.1].
Data availability No data was used for the research described in the article.
Declarations
Competing interests The authors declare that there is no conflict of interest.